OpenStudy (anonymous):

Help?? Darby bought a house for $445,000. She financed $356,175 of the purchase price with a 25-year, fixed-rate mortgage with a 7.15% interest rate. What is the total cost of the principal and interest after 25 years? a. $765,468.00 b. $956,364.00 c. $774,442.25 d. $967,577.25

OpenStudy (amistre64):

total sum of the payments of course. how much is a monthly payment?

OpenStudy (anonymous):

Is there a formula that goes with this?

OpenStudy (amistre64):

yes, its in your materials

OpenStudy (amistre64):

payment formula ...

OpenStudy (anonymous):

Would it be P = r(PV)/1-(1+r)^-n?

OpenStudy (anonymous):

Thanks. Hope im right. :)

OpenStudy (amistre64):

it not D, and your formula looks correct

OpenStudy (amistre64):

I never use that form tho, ive never liked it my form goes like this: \[Bk^n\frac{k^n-1}{k-1}=P\] we have our loan amount as B our compounding setup is k and n is the number of compounding periods

OpenStudy (amistre64):

356175k^(25*12)(k-1)/(k^(25*12)-1), k=1+.0715/12 http://www.wolframalpha.com/input/?i=356175k%5E%2825*12%29%28k-1%29%2F%28k%5E%2825*12%29-1%29%2C+k%3D1%2B.0715%2F12 then we simply add up all the payments made of 25 years

OpenStudy (amistre64):

If you take any random answer from a spammer, your bound to 'not get it right'

OpenStudy (anonymous):

Ok thanks! So I just take $2551.56 and multiply it by 25? Sorry, I really suck at math...

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

Wait. Thats not one of my answer choices... That was probably wrong, huh?

OpenStudy (amistre64):

25 years of payments, 12 a year 25*12*P should get us where we need to be

OpenStudy (amistre64):

it was almost right lolo

OpenStudy (anonymous):

Oh! Its A! Right?

OpenStudy (amistre64):

765... yeah, thats what my calcs get me

OpenStudy (anonymous):

Yay! Thank you!!!

OpenStudy (amistre64):

good luck